# TriSAM: Tri-Plane SAM for zero-shot cortical blood vessel segmentation in VEM images

### 25 Jan 2024 v1

@@ A.1 Failure Cases
Figure 11: The failure cases of the proposed method. Three prediction and ground truth pairs from BvEM-Macaque are shown. Results are generated from one initial seed located at the origin point.
A.2 Algorithm
The algorithm of TriSAM is shown in Algorithm 1.
Algorithm 1 TriSAM algorithm
[…]
Initialize the turning point tree $TurningPointTree=s$
for ${\textnormal{t}}_{i}\quad in\quad\operatorname{UnsampledLeafNode}(TurningPointTree)$ do
[…]
     if $\operatorname{Valid}(TurningPoints)$ then
         $TurningPointTree=\operatorname{LeafSplit}(TurningPointTree,{\textnormal{t}}_{i},TurningPoints)$triangleright Grow the turning point tree
     end if
end for
[…]
Refer to captionRefer to captionRefer to caption
A.3 Dataset statistics
The histogram of the blood vessels in the proposed dataset is shown in Figure 12.

### 09 Apr 2024 v2

@@ V-D Failure Cases
VI Conclusion
In this paper, we contribute the largest-to-date public benchmark, the BvEM dataset, for cortical blood vessel segmentation in 3D VEM images. We also developed a zero-shot blood vessel segmentation method, TriSAM, based on the powerful SAM model, offering an efficient and accurate approach for segmenting blood vessels in VEM images. With tri-plane selection, SAM-based tracking, and recursive seed sampling, our TriSAM effectively exploits the 3D blood vessel structure and attains superior performance compared with existing zero-shot and supervised technologies on BvEM across three species, marking a critical step towards unlocking the mysteries of neurovascular coupling and its implications for brain health and pathology. With the availability of the BvEM dataset and the TriSAM method, researchers are now equipped with valuable tools to drive breakthroughs in VEM-based cortical blood vessel segmentation and further our understanding of the brain’s intricate vascular network. By addressing a significant gap in the field of neuroimaging, we have laid the foundation for advancing the understanding of cerebral vasculature at the microscale and its intricate relationship with neural function.